Mark buys Call and Put options with different strike prices. Please read the following information.

Call option premium: $0.02/€

Put option premium: $0.01/€

Call option strike price: $1.15/€

Put option strike price: $1.05/€

Option contract size: €50,000

Break-even points for Mark are _______.

a. None of the above b. $1.02/€ and $1.18/€ c. $1.08/€ and $1.12/€ d. $1.06/€ and $1.16/€
Answer #1

Option B

Upper Breakeven Point = Strike Price of Long Call + Net Premium Paid=1.15+0.02+0.01=1.18

Lower Breakeven Point = Strike Price of Long Put - Net Premium Paid=1.05-0.02-0.01=1.02